QUESTION 9 During Year 1, El Paso Company had the following changes in account balances The Accumulated Depreciation account had a beginning balance of $62,500 and an ending balance of $87,500. The increase was due to depreciation expense. The Long-Term Notes Payable account had a beginning balance of $100,000 and an ending balance of $45,000. The decrease was due to repayment of debt. The Equipment Account had a beginning balance of $75,000 and an ending balance of $232,500. The increase was due to the purchase of other operational assets. The Long-Term Investments Account (Marketable Securities) had a beginning balance of $54,000 and an ending balance of $37,500 The decrease was due to the sale of investments at cost. . The Dividends Payable account had a beginning baance of $36,000 ad an ending balance of $30,000. There were $60,000 of dividends declared during the period. The Interest Payable account had a beginning balance of $6750 and an ending balance of $3750. The difference was due to the payment of interest. What is the net cash flow from financing activities? $66,000 inflow $121,000 outflow $55,000 inflow $55,000 outflowOn January 1, Year 1, Weller Company issued bonds with a $230,000 face value, a stated rate of interest of 10.00%, and a 10-year term to maturity. Weller uses the effective interest method to amortize bond discounts and premiums. The market rate of interest on the date of issuance was 8.00%. Interest is paid annually on December 31 Assuming Weller issued the bond for $248,240, what is the amount of interest expense that will be recognized during Year 3? (Round your intermediate calculations and final answer to the nearest whole dollar amount.) $23,000 $19,337 $19,859 $26,663On January 1, Year 1, Weller Company issued bonds with a $230,000 face value, a stated rate of interest of 10.00%, and a 10-year term to maturity. Weller uses the effective interest method to amortize bond discounts and premiums. The market rate of interest on the date of issuance was 8.00%. Interest is paid annually on December 31.
Answer #1

9)Correct option is "B"

Repayment of long term note payable [100000-45000] (55000)
Dividend paid [36000+60000-30000] (66000)
cash flow from financing activity (121000)

10)Interest paid = 230000*10% =23000

Date Interest paid Interest expense premium amortised carrying value at end
year 1 23000 248240*8%= 19859.2 3140.8 248240-3140.8= 245099.2
2 23000 19607.94   [245099.2*.08] 3392.06 245099.2-3392.06= 241707.14
3 23000 19336.57 3663.43 238043.71

Correct option is "B" - 19337